# Application of Battery Storage to Switching Predictive Control of Power Distribution Systems Including Road Heating*

### v1

@@ III-B Optimal Switching of Heating Cable
    In (9), $\Gamma_{\rm e}(x;t)$ [W/m] and $\Gamma_{\rm h}(x;t)$ [W/m] represent the distribution loss of the underground power distribution line and heating cable, respectively. Since the distribution loss of underground distribution line is considered for three phases, the distribution loss for a single-phase is three times larger and is given by
@@
  • The second term: The function $V_{\rm fluc}{[k]}$ [V] corresponds to the evaluation of the voltage fluctuation of the underground power distribution line in the time interval $[t,t+T_{\rm pred})$, and $w_{\rm fluc}$ is the weighting coefficient in the evaluation of the fluctuation. The function $V_{\rm fluc}[k]$ is defined by the maximum absolute values of the difference between the voltage amplitude and the reference value $v_{1}$ [V] for $N$ intervals:
@@ III-C Optimal Switching of PV Power Generation Equipment
In this subsection, we formulate the evaluation function that determines the destination of power electricity from the PV power generation equipment. Specifically, this function $J_{\rm pv}[k]$ is mathematically defined by
@@ III-D Optimal Discharging of Battery Storage
  • The first term: The function $V_{\rm batteryfluc}[k]$ [V] correspond to the evaluation of the voltage fluctuation of the underground distribution line in the time interval $[t,t+T_{\rm pred})$, and is defined in the same way as in the definition (10). Moreover, $w_{\rm batteryfluc}$ is the weighting coefficient of its evaluation.
@@ IV Numerical Verification
In this section, we verify the effectiveness of the proposed predictive switching control via numerical simulations considering two cases of typical power consumption, PV electric power generation, and weather conditions. These simulations suppose the winter season of Toyama Prefecture which is one of the heaviest snowfall areas in Japan.
@@ IV-A Simulation Setting
In this subsection, we provide the setting of the power distribution system considered in the numerical simulation. We consider a power distribution system consisting of the underground power distribution line and the heating cable with length $100$ m. The single-line diagram of the system is illustrated in Fig. 4. For the underground power distribution line, the reference voltage amplitude for p.u. value is set to $v_{1}=\frac{6600}{\sqrt{3}}$ [V]. Since we deal with only one of the three phases of the underground distribution line in this paper, we convert this reference value of line voltage  $\sqrt{3}v_{1}=6600$ [V] to a phase voltage. In the heating cable, the reference voltage amplitude for the p.u. value is set to $\frac{1}{a_{1}}v_{1}=200$ [V], where $a_{1}=\frac{33}{\sqrt{3}}$ is the transformer ratio. The load electrical equipment is installed at the horizontal position $x=25$ [m] from Bifurcation point 1, and the PV electric power generation equipment is installed at $x=50$ [m]. In addition, the battery storage is connected at $x=75$ [m]. The connecting position $x_{\rm load}$ [m] of the load of the heating cable is given by
@@
We describe the parameters used in the simulations of this section. The parameters related to the voltage distribution in the underground power distribution line and heating cable are shown in Tables I and II, respectively. For underground power distribution line and heating cable, the reference voltage amplitudes for the p.u. values are $v_{1}=6600$ [V] and $\frac{1}{a_{\rm A}}v_{1}=200$ [V], respectively. We assume that no reactive power is supplied at all position $x$ of the underground power distribution line and heating cable, i.e. $q_{\rm e}(x,t)=q_{\rm h}(x,t)=0$ [p.u.]. The parameters related to the temperature distribution of the heating cable, the thermal diffusion in the underground, and the snowmelting on the ground surface are shown in Table III. The parameters related to the evaluation function are summarized in Table IV. We set the prediction horizon to be $T_{\rm pred}=30$ [min] from the viewpoint of the prediction of temporal changes of snow melting against the occurrence of pedestrian falls and traffic congestion due to snow accumulation. Moreover, the minimum switching time is set to $T_{\rm mini}=10$ [min], considering the influence of frequent switching on the power distribution system.
@@
Voltage amplitude
of heating cable
@@ IV-B Time Series Data for Simulation
In this section, we describe the acquisition and processing of actual time series data on electricity and weather used in the simulations of this chapter.
IV-B1 Power consumption and PV electric power generation
The time series data on the power consumption of the residential loads are provided by Nagoya University Open Data for EMS Evaluation [16] for ten residential houses. As we consider only single phase of the three-phase underground distribution line, 1/3 of the time series data for 10 houses is used as the active power $p_{\rm e}(x,t)$ of the residential load.
We use time series data of PV electric power generation provided by Hokuriku Electric Power Transmission and Distribution Company [17]. We also use the value that is $\frac{1}{1000}$ of the data in order to fit our data to the scale of the underground power distribution line considered in this paper. Furthermore, as we consider only one phase among the three phases, we use 1/3 of the time series data that fits the scale of the distribution system considered in this paper as the effective power $p_{\rm e}(x,t)$ of the PV power generation equipment. On the other hand, when the power electricity is transmitted from the PV power generation equipment to the battery storage, we suppose that all the power electricity is transmitted. For this reason, we do not modify the scale of the data which is fit to the scale considered in this paper.
The power supplied by the PV electric power generation is considered to be more accurately predicted based on the day’s weather forecast, compared to the power consumption of the load. Thus, we use the time series data for PV electric power generation as predictive data. In general, it is not easy to predict these data exactly. For this reason, the predicted value of the data at the predictive horizon continues to use the value at the current time as a constant value in the computation of the evaluation function.
@@ IV-C1 Setting and time series data
In this subsection, we consider a numerical simulation for a sunny winter morning with no snowfall from A.M. 6:00 to 8:00. This implies that that A.M. 6:00 and 8:00 correspond to $t=0$ [min] and $t=120$ [min], respectively. For the time series data used in this case, we use January 19, 2016 for the residential load [16], January 19, 2020 for the PV electric power generation [17]. Fig. 5 shows the temporal variation of the power consumption of residential loads and PV electric power generation. We also use the time series data for atmospheric temperature [19], wind speed [19], and the average data from January 19, 1981 to 2009 for solar radiation [18]. The time series data on these quantities are shown in Fig. 6.
@@ IV-C2 Simulation result
Temporal variation of variables
We show the simulation results for Case 1 in Fig. 7. In Fig. 7 (a), the voltage amplitude $v_{\rm e}(x;t)$ [V] of the underground power distribution line varies as power consumption of the residential load and heating cable, and power supply from PV. The amplitude is reduced due to the power consumption of the residential load during A.M. 6:50 to 7:00. On the other hand, the amplitude increases with the PV electric power generation between A.M. 7:50 and 8:00. Fig. 7 (b) shows that the voltage amplitude $v_{\rm h}(x;t)$ [V] of the heating cable varies greatly as the switches change. When Switch 1 is turned On, the voltage amplitude decreases from $x=0$ [m] to $x=100$ [m]. On the other hand, the amplitude decreases from $x=100$ [m] to $x=0$ [m] during Switch 2 is turned On.
@@
(a) Voltage amplitude of underground distribution line
@@
Temporal variation of switching and discharging
@@
Comparison with and without battery storage
@@ IV-D1 Setting and time series data
In this subsection, we consider a numerical simulation for a winter evening from P.M. 4:00 to 6:00 with snowfall starting from P.M. 5:00. This implies that P.M. 4:00 and 6:00 correspond to $t=0$ [min] and $t=120$ [min], respectively. We use the time series data of the residential load [16] from February 7, 2016, and of the PV electric power generation [17] from February 7, 2020. The temporal variation of the power consumption of residential load and PV electric power generation is shown in Fig. 10. The time series data for atmospheric temperature and wind speed are from February 7, 2020 [19]. Moreover, we choose the average solar radiation data from February 7, 1981 to 2009 [18]. The snowfall time series data gives a hypothetical snowfall of $12$ mm per $1$ hour assumed for Toyama Prefecture. The time series data of physical quantities related to weather are shown in Fig. 11.
@@ IV-D2 Simulation result
Temporal variation of variables
The simulation results for Case 2 are shown in Fig. 12. We show the temporal and spatial variations of the voltage amplitude of the underground distribution line in Fig. 12 (a). In this figure, the voltage fluctuation is suppressed by the power supplied by the PV power generation equipment against the increase of power consumption by the residential load from P.M. 5:00 to 5:30. In Fig. 12 (b), the voltage amplitude of the heating cable decreases from $x=100$ [m] to $x=0$ [m] because Switch 2 is set to On at many time intervals.
@@
(a) Voltage amplitude of underground distribution line
@@
Temporal variation of switching and discharging
@@
We show the temporal variation of the control input  $\sigma_{{\rm pv}}(t)$ of the PV power generation equipment in Fig. 13 (b). In the time interval from P.M. 5:00 to 5:30, the equipment transmits power to the underground power distribution line. The reason for this is considered to be the increase in the power consumption of residential loads during this interval. Moreover, Fig. 13 (c) shows the temporal variation of the subscript $i$ satisfying $\sigma_{{\rm battery},i}(t)=1$, i.e., the destination of the power from the battery storage. When Switch 2 is set to On, the storage always transmits the power to the heating cable.
@@
Comparison with and without battery storage
@@ V Conclusions
In this paper, we have studied the application of a battery storage to the predictive switching control of a power distribution system including road heating proposed by the authors [9]. The introduction of a battery storage increases the number of transmission destinations of power electricity from the PV generation equipment, and improves the power flow in terms of optimization. The predictive switching control is designed for the PV power generation equipment and the battery storage, respectively. Its effectiveness is demonstrated by numerical simulations using the actual time-series data. As a conclusion, we have confirmed the effectiveness of the introduction of battery storage compared to the authors’ previous study [9].
As a future work, in order to improve the control performance of the switching predictive control, it is necessary to add constraints on the voltage amplitude and to consider the weights of the evaluation function according to the distribution system and its specific situation.

### v2

@@ III-B Optimal Switching of Heating Cable
    In (9), $\Gamma_{\rm e}(x;t)$ [W/m] and $\Gamma_{\rm h}(x;t)$ [W/m] represent the distribution loss of the underground power distribution line and heating cable, respectively. Since the distribution loss of underground power distribution line is considered for three phases, the distribution loss for a single-phase is three times larger and is given by
@@
  • The second term: The function $V_{\rm fluc}{[k]}$ [V] corresponds to the evaluation of the voltage fluctuation of the underground power distribution line in the time interval $[t,t+T_{\rm pred})$, and $w_{\rm fluc}$ is the weighting coefficient in the evaluation of the fluctuation. The function $V_{\rm fluc}[k]$ is defined by the maximum absolute values of the difference between the voltage amplitude and the reference value $v_{1}$ [V] for $N$ time intervals:
@@ III-C Optimal Switching of PV Power Generation Equipment
In this subsection, we formulate the evaluation function that determines the destination of power electricity from the PV power generation equipment. This function is mathematically defined by
@@ III-D Optimal Discharging of Battery Storage
  • The first term: The function $V_{\rm batteryfluc}[k]$ [V] correspond to the evaluation of the voltage fluctuation of the underground power distribution line in the time interval $[t,t+T_{\rm pred})$, and is defined in the same way as in the definition (12). Moreover, $w_{\rm batteryfluc}$ is the weighting coefficient of its evaluation.
@@ IV Numerical Verification
In this section, we verify the effectiveness of the proposed predictive switching control via numerical simulations considering two cases of typical power consumption, PV power generation, and weather conditions. These simulations suppose the winter season of Toyama Prefecture which is one of the heaviest snowfall areas in Japan. The simulation setting is given in Subsection  IV-A. In addition, the actual time series data for power consumption, PV power generation, and weather conditions used in this simulation are described in Subsection IV-B. Moreover, the simulation in Subsection IV-C considers the morning of a sunny winter day. Finally, in Subsection IV-D, we deal with a situation in which snowfall occurs, assuming an evening in winter. In particular, in both simulations, a comparison with the results of previous study [9] is given to illustrate the contribution of this paper.
@@ IV-A Simulation Setting
In this subsection, we provide the setting of the power distribution system considered in the numerical simulation. We consider a power distribution system consisting of the underground power distribution line and the heating cable with length $100$ m. The single-line diagram of the system is illustrated in Fig. 4. For the underground power distribution line, the reference voltage amplitude for p.u. value is set to $v_{1}=\frac{6600}{\sqrt{3}}$ [V]. Since we deal with only one of the three phases of the underground power distribution line in this paper, we convert this reference value of line voltage  $\sqrt{3}v_{1}=6600$ [V] to a phase voltage. In the heating cable, the reference voltage amplitude for the p.u. value is set to $\frac{1}{a_{1}}v_{1}=200$ [V], where $a_{1}=\frac{33}{\sqrt{3}}$ is the transformer ratio. The load electrical equipment is installed at the horizontal position $x=25$ [m] from Bifurcation point 1, and the PV power generation equipment is installed at $x=50$ [m]. In addition, the battery storage is connected at $x=75$ [m]. The connecting position $x_{\rm load}$ [m] of the load of the heating cable is given by
@@
We assume that no reactive power is supplied at all position $x$ of the underground power distribution line and heating cable, i.e. $q_{\rm e}(x,t)=q_{\rm h}(x,t)=0$ [p.u.].
@@
We describe the parameters used in the simulations of this section. The parameters related to the voltage distribution in the underground power distribution line and heating cable are shown in Tables I and II, respectively. The parameters related to the temperature distribution of the heating cable, the thermal diffusion in the underground, and the snow melting on the ground surface are shown in Table III. The parameters related to the evaluation function are summarized in Table IV. We set the prediction horizon to be $T_{\rm pred}=30$ [min] from the viewpoint of the prediction of temporal changes of snow melting against the occurrence of pedestrian falls and traffic congestion due to snow accumulation. Moreover, the minimum switching time is set to $T_{\rm mini}=10$ [min], considering the influence of frequent switching on the power distribution system.
@@
Voltage amplitude of heating cable
@@ IV-B Time Series Data for Simulation
In this subsection, we describe the time series data and processing of actual time series data on electricity and weather.
IV-B1 Power consumption and PV power generation
The time series data on the power consumption of the residential loads are provided by Nagoya University Open Data for EMS Evaluation [18] for ten residential houses. As we consider only single phase of the three-phase underground power distribution line, 1/3 of the time series data for 10 houses is used as the active power $p_{\rm e}(x,t)$ of the residential load.
We use time series data of PV power generation provided by Hokuriku Electric Power Transmission and Distribution Company [19]. We also use the value that is $\frac{1}{1000}$ of the data in order to fit our data to the scale of the underground power distribution line considered in this paper. Furthermore, as we consider only one phase among the three phases, we use 1/3 of the time series data that fits the scale of the distribution system considered in this paper as the effective power $p_{\rm e}(x,t)$ of the PV power generation equipment. On the other hand, when the power electricity is transmitted from the PV power generation equipment to the battery storage, we suppose that all the power electricity is transmitted. For this reason, we do not modify the scale of the data which is fit to the scale considered in this paper.
The power supplied by the PV power generation is considered to be more accurately predicted based on the day’s weather forecast, compared to the power consumption of the load. Thus, we use the time series data for PV power generation as predictive data. In general, it is not easy to predict these data exactly. For this reason, the predicted value of the data at the predictive horizon continues to use the value at the current time as a constant value in the computation of the evaluation function.
@@ IV-C1 Setting and time series data
In this subsection, we consider a numerical simulation for a sunny winter morning with no snowfall from A.M. 6:00 to 8:00. This implies that that A.M. 6:00 and 8:00 correspond to $t=0$ [min] and $t=120$ [min], respectively. For the time series data used in this case, we use January 19, 2016 for the residential load [18], January 19, 2020 for the PV power generation [19]. Fig. 5 shows the temporal variation of the power consumption of residential loads and PV power generation. We also use the time series data for atmospheric temperature [21], wind speed [21], and the average data from January 19, 1981 to 2009 for solar radiation [20]. The time series data on these quantities are shown in Fig. 6.
@@ IV-C2 Simulation result
(1) Temporal variation of variables
We show the simulation results for Case 1 in Fig. 7. In Fig. 7 (a), the voltage amplitude $v_{\rm e}(x;t)$ [V] of the underground power distribution line varies as power consumption of the residential load and heating cable, and power supply from PV. The amplitude is reduced due to the power consumption of the residential load during A.M. 6:50 to 7:00. On the other hand, the amplitude increases with the PV power generation between A.M. 7:50 and 8:00. Fig. 7 (b) shows that the voltage amplitude $v_{\rm h}(x;t)$ [V] of the heating cable varies greatly as the switches change. When Switch 1 is turned On, the voltage amplitude decreases from $x=0$ [m] to $x=100$ [m]. On the other hand, the amplitude decreases from $x=100$ [m] to $x=0$ [m] during Switch 2 is turned On.
@@
(a) Voltage amplitude of underground power distribution line
@@
(2) Temporal variation of switching and discharging
@@
(3) Comparison with and without battery storage
@@ IV-D1 Setting and time series data
In this subsection, we consider a numerical simulation for a winter evening from P.M. 4:00 to 6:00 with snowfall starting from P.M. 5:00. This implies that P.M. 4:00 and 6:00 correspond to $t=0$ [min] and $t=120$ [min], respectively. We use the time series data of the residential load [18] from February 7, 2016, and of the PV power generation [19] from February 7, 2020. The temporal variation of the power consumption of residential load and PV power generation is shown in Fig. 10. The time series data for atmospheric temperature and wind speed are from February 7, 2020 [21]. Moreover, we choose the average solar radiation data from February 7, 1981 to 2009 [20]. The snowfall time series data gives a hypothetical snowfall of $12$ mm per $1$ hour assumed for Toyama Prefecture. The time series data of the physical quantities related to weather are shown in Fig. 11.
@@ IV-D2 Simulation result
(1) Temporal variation of variables
The simulation results for Case 2 are shown in Fig. 12. We show the temporal and spatial variations of the voltage amplitude of the underground power distribution line in Fig. 12 (a). In this figure, the voltage fluctuation is suppressed by the power supplied by the PV power generation equipment against the increase of power consumption by the residential load from P.M. 5:00 to 5:30. In Fig. 12 (b), the voltage amplitude of the heating cable decreases from $x=100$ [m] to $x=0$ [m] because Switch 2 is set to On at many time intervals.
@@
(a) Voltage amplitude of underground power distribution line
@@
(2) Temporal variation of switching and discharging
@@
We show the temporal variation of the control input  $\sigma_{{\rm pv}}(t)$ of the PV power generation equipment in Fig. 13 (b). From P.M. 5:00 to 5:30, the equipment transmits power to the underground power distribution line. The reason for this is considered to be the increase in the power consumption of residential loads during this interval. Moreover, Fig. 13 (c) shows the temporal variation of the subscript $i$ satisfying $\sigma_{{\rm battery},i}(t)=1$, i.e., the destination of the power from the battery storage. When Switch 2 is set to On, the storage always transmits the power to the heating cable.
@@
(3) Comparison with and without battery storage
@@ V Conclusions
In this paper, we have studied the application of a battery storage to the predictive switching control of a power distribution system including road heating proposed by the authors [9]. The predictive switching control including the switching of the PV power generation equipment and the battery storage has been proposed as a main result. We have demonstrated the effectiveness by numerical simulations using the actual time series data. As a conclusion, we have confirmed the efficiency of the introduction of battery storage compared to the authors’ previous studies [9][11].
As a future work, to improve the control performance of the proposed control, it is necessary to add constraints on the voltage and current amplitudes, and to consider appropriate weights for the evaluation functions. The maximum hands-off control [22] can be introduced to reduce the number of switching events. In addition, we need to extend the proposed switching predictive control to the framework which also consider reactive power injection in order to suppress voltage phase delays and voltage fluctuations. A preliminary result has been obtained in the authors’ study [23].
This work was supported by Azbil Yamatake General Foundation and JSPS KAKENHI Grant Numbers 20K04552, 23K03910.